OpenStudy (anonymous):

Prove tan(θ / 2) = sin θ / (1 + cos θ) for θ in quadrant 1

OpenStudy (anonymous):

I substituted the tan(theta/2) with a half angle identity but I don't know where to go from there.

OpenStudy (anonymous):

\[\frac{ \sin \theta }{ 1+\cos \theta }\] \[\frac{ \ 2sin ( \theta /2 )* cos(\theta/2) }{ 2cos^2(\theta /2) }\] \[\frac{ \ 2sin ( \theta /2 )* cos(\theta/2) }{ 2cos(\theta /2)*cos(\theta /2) }\] \[\frac{ \sin \theta/2 }{ \cos \theta/2 }\] \[ tan (\theta/2) \]

OpenStudy (anonymous):

Q.E.D!

OpenStudy (anonymous):

how did you get from sinθ/1+cosθ to [2sin(θ/2)∗cos(θ/2)]/[2cos2(θ/2)]

OpenStudy (anonymous):

do you know the formulae for sin( A+B) and Cos(A+B) ??

OpenStudy (anonymous):

yeah: sin (a+b)=sinacosb+cosasinb and cos(a+b)= cosacosb-sinasinb

OpenStudy (anonymous):

But how does that relate? @SandeepReddy

OpenStudy (anonymous):

we have formulae: sin(x+y)=sinx.cosy+cosx.siny sin(x−y)=sinx.cosy−cosx.siny cos(x+y)=cosx.cosy−sinx.siny cos(x−y)=cosx.cosy+sinx.siny tan(x+y)=tanx+tany1−tanx.tany tan(x−y)=tanx−tany1+tanx.tany remember these so sin(x+x) = sin(2x) = 2sinx.cosx and cos(x+x) = cos^2x - sin^2x = 2cos^2x - 1

OpenStudy (anonymous):

i used the last two formulae

OpenStudy (anonymous):

\[ sin \theta = 2 sin\theta/2 * cos \theta/2 \]

OpenStudy (anonymous):

as we have taken 2x = theta/ then x becomes theta/2 right? same applies for cos2x

OpenStudy (anonymous):

do you know the half angle formulas for: \[\sin \frac{ \theta }{ 2 }\] and \[\cos \frac{ \theta }{ 2 }\] ? if yes, then you can prove it by going back to the formula of tangent. \[\tan \theta =\frac{ \sin \theta }{ \cos \theta }\] \[\tan \frac{ \theta }{2 }=\frac{\sin \frac{ \theta }{2 } }{ \cos \frac{ \theta }{ 2 } }\]

OpenStudy (anonymous):

substitute the half-angle formula of sine and cosine. \[\tan \frac{ \theta }{ 2 }=\frac{ \sqrt{\frac{ 1-\cos \theta }{ 2 }} }{\sqrt{\frac{ \cos \theta+1 }{ 2 }} }\] cancel out two. \[\tan \frac{ \theta }{ 2 }=\frac{ \sqrt{1-\cos \theta} }{ \sqrt{1+\cos \theta} }\] rationalize the denominator.

OpenStudy (anonymous):

\[\tan \frac{ \theta }{ 2 }=\frac{ \sqrt{1-\cos ^{2}\theta} }{ 1+cos }\] \[\tan \frac{ \theta }{ 2 }=\frac{ \sqrt{\sin ^{2}\theta} }{ 1+\cos \theta }\] \[\tan \frac{ \theta }{ 2 }=\frac{ \sin \theta }{ 1+\cos \theta }\] PROVED. :)

OpenStudy (anonymous):

@ECEstudent9405 how did you get to tanθ/2=√1−cos^2θ/1+cosθ?

OpenStudy (anonymous):

oh ok-ok. sorry. let's just focus on the numerator ok? when you rationalize it, it will be: \[(\sqrt{1-\cos \theta})(\sqrt{1+\cos \theta})\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

o i get it! Thanks :)

OpenStudy (anonymous):

wow. You're good. :) no problem.